OpenStudy (anonymous):

Show how changing the position of the parentheses affects the solution for the expression: 40 divided by 5 + 5 * (2^2 -1) Show all of your work

hero (hero):

Have you tried changing the position of the parentheses? For instance, would 5 + (5 * 2^2) -1 produce the same result as 5 + 5 * (2^2 -1) ?

OpenStudy (anonymous):

Thank you

hero (hero):

Wait, what I posted isn't the solution. I was just asking you if you had already tried switching the position of the parentheses.

OpenStudy (anonymous):

oohh

OpenStudy (anonymous):

idk this is my last question in my test can u just give me the answer

hero (hero):

Well, at least you're being honest.

hero (hero):

Yes, the result changes because... 5 + 5 * (2^2 -1) = 5 + 5*(4 - 1) = 5 + 5 * (3) = 5 + 15 = 20 5 + (5 *2^2) - 1 = 5 + (5*4) - 1 = 5 + 20 - 1 = 25 - 1 = 24

OpenStudy (anonymous):

thank you